A firm is considering a group of 3 contingent projects. Each project is 5 years long and the firm has a cost of capital of 9%. The firm is certain of the cash flows for project A and project B. However, the firm is not sure of the annual cash flow for project C. Using the information, find the minimum annual cash flow for project C (year 1 to year 5) for the firm to accept all 3 projects?

A B C
YEAR 0 CASH FLOW -$100,000.00 -$150,000.00 -$275,000.00
YEAR 1 CASH FLOW $25,000.00 $35,000.00 $???
YEAR 2 CASH FLOW $25,000.00 $35,000.00 $???
YEAR 3 CASH FLOW $25,000.00 $35,000.00 $???
YEAR 4 CASH FLOW $25,000.00 $35,000.00 $???
YEAR5 5 CASH FLOW $25,000.00 $35,000.00 $???
Answer #1

Sol. For calculating the annual cashflows for the Project C, we will use annuity formula:

Present value of annuity = C * [{1−(1+i)-n } / i ] Where, C = Annual Cash flow

275,000 = C * [1 - {(1+0.09)-5 } / 0.09 ] PV of annuity = Amount of outflow at year O

275,000 = C * [ {1 - 0.6499} / 0.09] i = cost of capital

275,000 = C * [0.3501/0.09] n = Period of incoming cashflow

C = 275000 / 3.89 = $70,694

Thus, the firms needs minimum $70,694 as inflows for all five years, in order to accept Project C